OpenStudy (anonymous):

At the beginning of the year 2000, Bob put $100 in a savings account. The bank pays Bob 5% interest on his total savings at the end of each year including all interest added to the account. Assume the interest rate continues and Bob does not deposit any additional money in the account. How much will he have in his savings account, to the nearest cent, after 5 interest payments? a. $150.00 c. $127.63 b. $155.55 d. $500

OpenStudy (texaschic101):

100(.05) = 5....100 + 5 = 105 (year 1) 105(.05) = 5.25...105 + 5.25 = 110.25 (year 2) 110.25(.05) = 5.51...110.25 + 5.51 = 115.76 (year 3) 115.76(.05) = 5.79...115.76 + 5.79 = 121.55 (year 4) 121.55(.05) = 6.08...121.55 + 6.08 = 127.63 (year 5) There is probably a formula for this...I just did it this way. Answer is C